19 September 2013: Thursday

LDAP:

  • Updates:
    • HR confirmed the match from yesterday and the merge was performed.
    • HR reported that the Job record had been updated, but it didn’t make the new feed file.  We’ll see if it does today.
    • One former student from Banner.  Added the LDAP entry.
  • Changes:
    • One VPN update performed.

Backups:

  • Savegroups:
    • One windows client decided to ignore backups last night.  Administrators notified.
    • The CEMS system is still being worked on.
  • Versions:
    • Upgraded the NW8.0sp2 repository to hold 8.0.2.3 (released on 9/16).
  • Tapes:
    • Re-labeling the tapes that came back from the Vault yesterday.
  • Test System:
    • 45 day install expired, rebuilding.
  • New Hardware:
    • Researching NFS Performance measurements/tuning tools.

Meetings:

  • Backup system configuration consultation:

 

Tags:

Comments are closed.